OpenStudy (anonymous):

x^2 + (x+20)^2= 1200

OpenStudy (anonymous):

\[x^2+x^2+40x+400=1200\] is a start then \[2x^2+40x=800\] or if you divide by 2 \[x^2+20x=400\]

OpenStudy (anonymous):

you know how to solve this?

OpenStudy (anonymous):

no and frustration is rising.

OpenStudy (anonymous):

^ ahahaha

OpenStudy (anonymous):

quadratic equation you speak it?

OpenStudy (anonymous):

Order of operations. First square what is in the parenthesis.

OpenStudy (anonymous):

\((x+20)(x+20)\) use FOIL.